# ECM: A Unified Electronic Circuit Model for Explaining the Emergence of In-Context Learning and Chain-of-Thought in Large Language Model

### v1

@@ 2.4 Model Performance as Output Power of Bulb
Here, output power represents the model’s effective performance, with higher power indicating better task execution, akin to a brighter bulb. The equation highlights that maximizing performance requires enhancing the basic voltage of power supply ($\mathcal{E}_{\text{model}}$) of LLM, minimizing reasoning resistor ($R_{\text{CoT}}$) in CoT, and optimizing the extra voltage ($\mathcal{E}_{\text{ICL}}$) provided by ICL. This synergy explains and predicts the model’s performance on complex tasks quantitatively.
Figure 2: The Experimental Results of Theoretical Model Verification on ECM.
a,
An analysis of ECM incorporating the CoT approach reveals a scatter plot illustrating the relationship between model accuracy and theoretical power with 0.8843 Spearman coefficient. b. Six similarity metrics are employed to evaluate whether context demonstrations can be interpreted as a semantic magnetic field. The results indicate that the projection length achieving the highest coefficient of 0.9106 between practical accuracy and theoretical power. c, The integrated ECM framework combines ICL and CoT. Across three retrieval strategies for ICL demonstrations, a consistent linear correlation is observed. d, Experiments on various LLMs under different power supply voltages, representing their capacities, reveal a strong linear correlation between theoretical power estimates and observed accuracy. e, Broader analyses across diverse tasks, including multi-hop knowledge reasoning, medical probing, and robotic planning, demonstrated robustness.
3 Theoretical Model Verification
3.1 CoT Meets Ohm’s Law
First, we verify whether CoT adheres to Ohm’s law. Since the electric current $I_{model}$ in Eq. 2 is not directly observable, we instead examine whether $P_{out}$ is linearly related to true accuracy, thereby indirectly validating CoT Ohm’s law. Specifically, to control variables, we remove the ICL and let the model perform zero-shot reasoning step-by-step following the instructions, which satisfies: $P_{\text{out}}=\frac{\mathcal{E}_{\text{model}}^{2}R_{0}}{(R_{\text{CoT}}+R_{0})^{2}}.$ The strong Spearman correlation coefficient between theoretical power ($P_{\text{out}}$) and observed accuracy (Acc), shown in Figure 2a, is 0.8843, with a $p$-value below 0.01. This strong correlation suggests difficulty of CoT reasoning actually exhibits resistor-like behavior, making it a reliable predictor of model performance.
3.2 ICL Meets Faraday’s Law
We conceptualize contextual demonstrations in ICL as a semantic magnetic field governed by Faraday’s Law during reasoning. Under this framework, semantic influence and proximity are quantified by projection length, as illustrated in the lower right of Figure 2b. To validate this hypothesis, we evaluate various metrics for contextual proximity between demonstrations and user requests. Our analysis (Figure 2b) shows that the Spearman correlation between the calculated power $P_{\text{model}}$ in Eq. 3 and observed spearman coefficient and R2 merely for projection length exceeds 0.85. This indicates that the geometric alignment of the demonstrations with the task request is a strong predictor of model performance, which further prove the existance of Faraday’s Law of ICL.
3.3 ICL and CoT Satisfy a Unified ECM
In an unprecedented integration, we unify ICL and CoT within a single ECM, offering a comprehensive framework to explain their interaction and collective impact on model performance. We rigorously test this unified model using three distinct retrieval strategies for ICL demonstrations, all of which yields a consistent linear distribution, as illustrated in Figure 2c. This consistent alignment, with a Spearman correlation coefficient of 0.8979 and a $p$-value approaching zero between calculated power in Eq. 3 and practical accuracy, provides strong empirical evidence for the robustness of our theoretical model. These results confirm that both ICL and CoT not only can be calculated independently but are inherently compatible within the same theoretical structure, offering a powerful new way to describe and predict the behavior of LLM systems.
3.4 Extension Experiments on Various LLMs and Tasks.
To demonstrate the generality of our theoretical ECM, we conduct extension experiments across LLMs configured with varying power supply voltages to represent different capacities. Figure 2d reveals a consistent linear correlation over 0.8348 between theoretical output power estimates and empirical accuracy, confirming the ECM’s robustness in predicting performance. Beyond traditional mathematical reasoning tasks, we test ECM on diverse real-world problems, including multi-hop knowledge reasoning [44], medical probing [7], and robotic planning [34]. As shown in Figure 2e, ECM achieves Spearman coefficients exceeding 0.86 and R2 exceeding 0.77 across all tasks, underscoring its versatility in both cognitive and practical applications.
4 Explanation based on ECM
In this section, we mainly pay attention to how to use ECM to explain the effectiveness of existing prompting methods.
4.1 Explanation for ICL Strategies
Enhanced embeddings facilitate accurate semantic magnetic field strength estimation and retrieval.
Recent advancements in embedding models have shown that improved text representations directly enhance LLM’s efficacy in the retrieval process [21]. This improvement is tied to the precise estimation of semantic magnetic field strength. As shown in Figure 3a, we employ three models—BERT [9], RoBERTa [22], and BGE [4]—to represent and retrieve demonstrations in ICL. Among these, BGE achieves the highest degree of alignment, with a Spearman correlation coefficient of 0.9106 between theoretical power and empirical accuracy, significantly surpassing the other models. This result highlights superior embedding ability attributed to the accurate semantic magnetic field estimation.
Figure 3: The Experimental Explanation for Existing ICL Strategies.
a,
The relationship between theoretical power and actual accuracy for various encoders is assessed. BGE achieves the highest Spearman correlation coefficient of 0.9106. b, A comparison of the Vanilla Encoder and the Meta-Recognition Representation reveals the latter’s ability to enhance the linear correlation between theoretical power and practical accuracy, from 0.9130 to 0.9337. c, The distribution of magnetic field strengths in generated demonstrations (with CDF representing cumulative distribution function and CCDF representing complementary cumulative distribution function). Further, we demonstrates the effectiveness of different Synthetic-CoT with degree of different degree of semantic field strength. d, The effectiveness of Synthetic-CoT is analyzed across three conditions: Positive Synthetic-CoT (demonstrations with positive extra voltage from ICL based on positive magnetic fields), Zero-shot CoT (no demonstrations), and Negative Synthetic-CoT (demonstrations with negative extra voltage from ICL based on negative magnetic fields). The left scatter plot displays the performance distribution of these conditions on the power-accuracy landscape. The right bar chart highlights factors contributing to Synthetic-CoT’s optimization of extra voltage provided by ICL.
Meta Recognition Enhance Magnetic Field Strength Estimation Modeling.
The study of Didolkar et al. [10] introduces Meta-Recognition, a method leveraging GPT-4 [1] to decompose complex problems into modular sub-capabilities for ICL representation, improving LLMs’ performance. We attribute its effectiveness to the enhanced sample representation modeling, which enhances ICL voltage and subsequently boosts output power and accuracy. As illustrated in Figure 3b, Meta-Recognition significantly increases the correlation between theoretical power and practical accuracy, improving from 0.9130 to 9337. It underscores that representation based on sub-capability decomposition effectively enhances the precision of magnetic field strength modeling.
Synthetic-CoT predominantly generates samples with positive magnetic fields, enhancing model performance.
Synthetic-CoT, a widely used approach in ICL, enables LLMs to autonomously generate ICL demonstrations, thereby improving performance [33]. As illustrated in Figure 3d, all ICL demonstrations produced by Synthetic-CoT exhibit positive semantic magnetic field contributions, which are strongly correlated with enhanced model accuracy. This effect is likely due to the positive voltage induced by the magnetic field, which optimizes circuit functionality. As shown in Figure 3d, the larger ICL magnetic field strength corresponds to the larger power and the higher accuracy. To further investigate the impact of magnetic field polarity on model performance, we analyzed demonstrations with negative magnetic field. Including such examples led to a 21.47% reduction in accuracy and a substantial decline in theoretical power output (Figure 3d). These findings underscore the critical role of positive magnetic field demonstrations in the superior performance achieved by Synthetic-CoT.
Figure 4: The Experimental Explanation for Existing CoT Strategies.
a,
The distributions of accuracy, power, and resistance for Zero-shot without CoT and Zero-shot with CoT are presented. The average resistance for Zero-shot without CoT is 10.93, significantly higher than the 1.50 observed for Zero-shot with CoT. b, Mechanistic analysis of Tool Usage (TU) and Program-of-Thought (PoT) within the ECM. Initially, TU, modeled on the original CoT, had non-zero computational resistance ($R_{C}$), limiting power-accuracy correlation. Assuming $R_{C}=0$ achieves a 10% improvement in the Spearman coefficient. PoT, representing TU with structured code, further enhanced performance by keeping $R_{C}=0$ and reducing planning resistance $R_{P}$, achieving another 10% improvement in power-accuracy correlation. c, A circuit-based rationale to explain how to utilize self-consistency and coverage to enhance performance. Self-consistency reduces total resistance by processing multiple reasoning resistors in parallel (sampling $n$ solutions and selecting the output with the highest consensus). In contrast, coverage minimizes resistance by increasing sample size with negligible resistance of voting resistor (sampling $n$ instances and passing the task if any succeeds).
4.2 Explanation for CoT Strategies
Direct Answering Introduces Greater Logical Resistor Compared to CoT
Recent studies suggest that generating direct answers without intermediary reasoning steps increases cognitive load on the model [13, 17]. We hypothesize that the absence of explicit reasoning pathways introduces significant resistance of “logical resistor,” leading to weaker performance compared to Zero-shot CoT. To test this, we analyzed performance and power correlations for direct answering and Zero-shot CoT on a power-accuracy graph, finding a strong linear correlation of 0.9131 ($p$-value $<0.01$). The average resistance for direct answering is calculated as 10.93, substantially higher than the 1.50 observed for Zero-shot CoT. This discrepancy corresponds to a performance reduction of over 30% for direct answering. These findings indicate that the lack of a CoT framework imposes additional greater logical resistor, hindering the model’s reasoning capabilities and resulting in diminished performance.
Tool Usage optimizes computational resistor, and PoT further reduces planning resistor, improving model performance.
Chen et al. [5] propose that Program of Thought (PoT) and Tool Usage (TU) enhance computational reasoning capabilities to a theoretical maximum, effectively minimizing reasoning complexity to zero, akin to eliminating the resistance to zero. Inspired by this, we investigate the correlation between CoT and TU models under this "zero-resistance" assumption. Our analysis revealed that modeling TU as a circuit with zero resistance of computational resistor increased the Power-accuracy Spearman correlation by over 10% compared to the original model, as illustrated in Figure 4b, which validates our hypothesis.
Furthermore, we hypothesize that PoT, leveraging programming languages, reduces planning resistance. As shown in Figure 4b, modeling PoT as a circuit with lower planning resistance achieved a correlation score exceeding 0.9423 between theoretical power and accuracy—significantly higher than TU. This finding highlights the superior effectiveness of PoT in improving model performance.
Self-consistency and Inference Scaling Law via Parallel Resistor in Circuits
Self-consistency, as demonstrated by Wang et al. [40], utilizes multiple reasoning paths simultaneously, and determines final decision through majority voting, enhancing performance robustness and accuracy. Here, we interpret its effectiveness using a circuit model, which introduce parallelization of original resistors ($R_{\text{CoT}}$), significantly reducing the total system resistance. As shown in Figure 4c, the practical accuracy of the self-consistency strategy aligns closely with the theoretical output power curve, validating the correctness of our proposed model. Recent findings on the Inference Scaling Law [43, 28] further demonstrate that increasing the number of parallel samples, $n$, substantially improves accuracy. As $n$ grows, the resistance value of CoT resistor approaches zero, optimizing the reasoning process. Further, when coverage introduces an absolutely correct answer verification process with zero resistance of verification resistor, exceeding the gains achieved by self-consistency alone. The alignment between theoretical predictions and empirical results, also evident in Figure 4c, reinforces the validity of this model. These results highlight the central role of self-consistency and inference scaling in enhancing the performance of circuits.
Figure 5: The Experimental Exploration for ICL Strategies.
a,
The relationship between the introduction of demonstrations and model performance. Using Reverse ICL + CoT, which introduces demonstrations with negative magnetic fields, impairs the model’s reasoning ability and results in performance falling below that of Zero-shot CoT. In contrast, Retrieval-ICL + CoT, incorporating demonstrations with positive magnetic fields, significantly improves performance. b, The upper part of the figure contrasts Diversity-oriented Static ICL (DS-ICL), where demonstrations are retrieved for the whole task, and Similarity-oriented Dynamic ICL (SD-ICL), where demonstrations are retrieved for each request. In SD-ICL, demonstrations are more aligned with test requests, creating a stronger semantic magnetic field. The lower part highlights Diversity-oriented Dynamic ICL (DD-ICL), where increasing sampling range enhances the diversity but decreases the magnetic field strength of retrieved demonstrations, resulting in a more diverse but weaker strength, leading to less accuracy. c, The left figure illustrates the demo-optimization process, a methodology designed to refine the demonstrations used in ICL. The right figure shows that demo-optimization outperforms traditional prompt optimization by improving theoretical power.
5 Exploration based on ECM
In this section, we primarily focus on how to use ECM to explore phenomena that have never been well explained and more novel prompting methods.
5.1 ICL Exploration
When Few-shot CoT Performs Worse than Zero-shot CoT?
Despite the growing importance of ICL and CoT, a critical question remains: under what conditions is ICL essential, and when does zero-shot CoT suffice? To investigate this, we systematically examine the interaction between these two strategies. As shown in Figure 5a, our findings reveal a phenomenon we term “reverse ICL," where demonstrations associated with a negative semantic magnetic field significantly impair the output power of circuits, thereby decreasing the model’s reasoning ability and resulting in worse performance compared to zero-shot CoT. This suggests that not all demonstrations contribute positively to the learning process. In contrast, demonstrations with a positive semantic magnetic field lead to an enhancement in performance. These results highlight the critical importance of selecting demonstrations based on their associated semantic magnetic field, which plays a crucial role in determining the success of ICL.
The influence of diversity in ICL retrieval: a static vs. dynamic perspective
In ICL retrieval, two paradigms are commonly studied: static ICL, where the set of demonstrations remains fixed during testing, and dynamic ICL, where demonstrations are selected in real-time based on the specific test sample. Previous studies [19] reveal that semantic diversity in static ICL enhances performance by exposing the model to a broader range of contextual cues. Conversely, in dynamic ICL, diversity does not provide the same benefits and may even hinder performance [31]. This raises the question: Why does diversity yield such contrasting outcomes in these retrieval modes? To this end, we hypothesize that the benefits of diversity in static ICL can be likened to a magnetic field. Diverse demonstrations create a robust and expansive field, offering the model access to varied directional signals as test samples change. This broad-spectrum magnetic field stabilizes the model’s decision-making process, improving its generalizability and overall performance. As illustrated in Figure 5b, static ICL benefits from this diversity, which amplifies the model’s predictive power compared to random selection strategies. In dynamic ICL retrieval, however, the focus shifts toward selecting demonstrations closely aligned with the test sample. This approach optimizes the magnetic field for stronger, more targeted signals in the relevant direction. In this context, diversity becomes minimally critical; the alignment and intensity of retrieved demonstrations take precedence. As shown lower panel in Figure 5b, diverse but less aligned demonstrations enable less accuracy.
Demo-Optimization: a novel approach to enhance demonstration efficacy in ICL
Recent advancements in prompt-based learning leverage feedback from outputs as model gradients, enabling large models to autonomously refine their prompts for more efficient learning [45]. Expanding on this foundation, we propose demo-optimization, a methodology tailored to optimize the demonstrations used in ICL. Demo-optimization focuses on balancing the relative effectiveness between user requests and demonstrations (See Method for implementation details). As illustrated in Figure 5c, demo-optimization significantly outperforms traditional prompt optimization approaches, which primarily refine instructions. This method not only improves accuracy but also theoretically strengthens reasoning capabilities. Moreover, it addresses the persistent underfitting problem in prior ICL methods, offering a more precise and scalable strategy for enhancing demonstration-driven learning.
Figure 6: Experimental Evaluation of Chain-of-Thought Strategies and Power-Optimized Performance on Complex Tasks
a,

### v2

@@ 2.4 Model Performance as Output Power of Bulb
Here, output power represents the model’s effective performance, with higher power indicating better task execution, akin to a brighter bulb. The equation highlights that maximizing performance requires enhancing the basic voltage of power supply ($\mathcal{E}_{\text{model}}$) of LLM, minimizing reasoning resistor ($R_{\text{ITR}}$) in ITR, and optimizing the extra voltage ($\mathcal{E}_{\text{ITL}}$) provided by ITL. This synergy explains and predicts the model’s performance on complex tasks quantitatively.
Figure 3: Verification of the ECP via zero‐shot reasoning.
a, Log-linear relationship between the classical inference-time scaling law (log token length versus task accuracy) and empirical performance of 4 LLMs, GPT-3.5 Turbo, GPT-4o-mini, GPT-4o and Qwen2.5-7B-Instruct, reveals a significant negative correlation (Spearman $\rho\in[-0.5664,-0.4254]$,$P<0.01$). b, Observed zero-shot accuracy plotted against ECP-predicted computational power for 9 LLMs shows a strong positive correlation (Spearman $\rho>0.71$, $P<0.01$), confirming the ECP’s predictive validity. c, Comparative evaluation of 4 semantic-similarity metrics used to estimate field strength: projection length, Manhattan ($L_{1}$) distance, Euclidean ($L_{2}$) distance and cosine similarity, identifies projection length as the most accurate predictor. More complete results are shown in Fig. 9 in Appendix.
3 Results
3.1 Accurate Performance Prediction
3.1.1 ITR Meets Ohm’s Law to Accurately Predict Zero-shot Performance
To investigate whether ITR obeys Ohm’s law, we eliminate all inference-time learning effects (i.e., we exclude any in-context demonstrations in the prompts), so that $\mathcal{E}_{\mathrm{ITL}}=0$. Since the model current $I_{\mathrm{model}}$ in Eq. 2 be measured directly, we instead quantify the output power to assess ITR’s adherence to Ohm’s law, which can be calculated as follows:
$Acc\propto P_{\mathrm{out}}=\frac{\mathcal{E}_{\mathrm{model}}^{2}\,R_{0}}{(R_{\mathrm{ITR}}+R_{0})^{2}}.$ (4)
As shown in Fig. 3a, the traditional Inference Time Scaling Law, predicting performance from token length alone, fails to predict reasoning performance accurately (Spearman’s $\rho\approx-0.5$ and Pearson’s $r<0.53$). In contrast, as illustrated in Fig. 3b, the output power of our ECP correlates strongly and positively with actual accuracy across every model tested ($\rho,r>0.7$). Our findings reveal that reasoning difficulty functions analogously to electrical resistance: as cognitive “resistance” increases, the “current” of information flow, and the relevant “power” of reasoning output diminishes, echoing the dynamics of Ohm’s law.
3.1.2 ITL Meets Faraday’s Law to Accurately Predict Few-shot Performance
Next, we introduce a novel conceptual framework for ITL, drawing an analogy to Faraday’s law: contextual demonstrations can be interpreted as generating a “semantic magnetic field” that influences task performance. To probe this analogy, we compare four proximity metrics: projection length, a baseline with no distance modelling, Euclidean distance, and cosine distance, quantifying the relationship between demonstrations and task prompts. As shown in Fig. 3c, projection length exhibits the strongest correlation with task accuracy ($\rho=0.8251$, $r=0.8481$, $R^{2}=0.6095$), outperforming both the No Distance baseline ($\rho=0.7913$, $r=0.8320$, $R^{2}=0.5555$) and conventional metrics such as Euclidean distance ($\rho=0.7597$, $r=0.8170$, $R^{2}=0.6017$) and cosine distance ($\rho=0.8151$, $r=0.8217$, $R^{2}=0.5190$). Strikingly, the latter two, even though widely used for demonstration retrieval, perform worse than having no distance modelling at all. These findings reveal that projection length between demonstrations and task requests is another key to effectively model ITL effectiveness.
Figure 4: Interpreting for Inference‐Time Learning Strategies of ECP.
a, ECP‐predicted theoretical power correlates positively with accuracy across three semantic encoders (BERT, RoBERTa, BGE) over 350 tasks. b, Meta‐recognition embeddings yield higher mean theoretical power (73.6 vs 66.9) and accuracy (54.8% vs 49.9%) than vanilla encoders, with stronger Spearman correlations (0.75 vs 0.61, $P<0.01$), which using RoBERTa as semantic encoders. c, Cumulative distributions (CCDF, CDF) and 95% confidence ellipses for three ITR configurations (Acc = 0.442, 0.376, 0.352) illustrate that larger induced semantic fields correspond to higher theoretical power and observed accuracy.
3.2 Modular Performance Optimization
3.2.1 Interpreting Existing Optimization Strategies by Magnetic Field Adjustment
Enhanced semantic encoders sharpen magnetic-field estimates and thereby improve retrieval accuracy.
Recent progress in neural semantic encoders has markedly improved the performance of ITL tasks [26]. To understand the underlying reason, we compared three prominent encoders (BERT [9], RoBERTa [27], and the more recent BGE [3]), by correlating power-accuracy metrics (Fig. 4a). BGE exhibited the strongest agreement with ground truth (Spearman’s $\rho=0.83$, Pearson’s $r=0.85$, $R^{2}=0.61$), surpassing its predecessors and indicating that higher-quality embeddings generate more coherent semantic fields and more reliable performance forecasts.
Meta-recognition representations further augment magnetic-field modelling.
[…]
Synthetic-CoT consistently consistently produces positive magnetic fields.
Synthetic-CoT requires LLMs to generate demonstrations at inference [37]. Stronger fields correlate positively with retrieval accuracy, as shown by complementary cumulative distributions and 95% confidence ellipses in power–accuracy space (Fig. 4c). High-field Synthetic-CoT achieved an accuracy of 0.442, outperforming its low-field counterpart (0.376) and the zero-shot baseline (0.352). These converging results identify field magnitude as a key determinant of ITL performance and outline scalable principles for tuning magnetic fields through representation learning and prompt engineering.
Figure 5: Interpreting for Inference‐Time Learning Strategies of ECP.
a, ITR reconfigures the model’s reasoning circuit: relative to Rationale, removing the rationale (w/o Rationale) increases effective resistance (3.97 $\Omega$ vs 4.40 $\Omega$), lowers theoretical power (62.1 vs 56.7) and improves accuracy (41.7% vs 33.6%), as shown by the contour shifts. b, Applying natural language (NL), tool usage (TU) or program-of-thought (PoT) to instantiate the “superconducting” phenomenon ($R_{c}=0$) surpasses NL for better Power-Accuracy correlations, in line with findings on externalized reasoning. c, Circuit analogies for two sampling strategies. Self-consistency sampling comprises parallel reasoning paths merging through an aggregation resistance $R_{s}$, denoting consensus cost; coverage sampling sets ($R_{s}=0$), allowing independent paths for greater diversity. d, Empirical trends follow a logarithmic relation with sample count. e, Adopting independent rather than mixture sample counts motivates replacing linear models with logarithmic ones, markedly improving predictions.
3.2.2 Interpreting Existing Optimization Strategies through Resistor Reorganization
Direct answering increases logical resistance relative to zero-shot ITR.
Prompting LLMs to produce answers directly, without exposing their intermediate reasoning, imposes greater cognitive demands and reduces task performance [43, 21], which we attribute it to imposing additional “logical resistance” on the inference pathway. As illustrated in Fig. 5a, we found that direct answering lowers the mean theoretical power requirement (62.10 $\rightarrow$ 56.65)but simultaneously raises effective logical resistance from 3.97 to 4.40, driven by 1.1×, 1.5× and 1.6× increases in planning, calculation and local‐operation resistances, respectively. Accuracy consequently declines from 41.7% to 33.6%, indicating that explicit chains of thought enhance energy efficiency and steer the model towards correct solutions.
Tool usage and program-of-thought minimise calculation and planning loads.
Following Chen et al. [4], representing reasoning complexity as an ohmic load, we tested whether explicit tool calls (tool usage, TU) and step-wise code generation (program-of-thought, PoT) approach a “superconducting” limit. As shown in Fig. 5b, setting the computational resistor to zero improved the power–accuracy fit by $8\%$ in Pearson’s $r$ and by $10\%$, supporting the zero‐resistance “superconducting” hypothesis for TU. Further reducing the planning resistor for PoT produced an even tighter correspondence ($r>0.95$ and $R^{2}>0.89$), demonstrating PoT’s superior efficiency in converting computational throughput into correct answers.
Self-consistency and inference scaling act as parallel resistors.
As shown in Fig. 5c, executing $n$ concurrent reasoning chains and selecting the majority answer, also known as Self-consistency [41], can be modelled as $n$ parallel resistors of value $R_{\text{ITR}}$ which reduce the effective resistance to $R_{\text{ITR}}/n$; a verification step of resistance $R_{S}$ sets a lower bound (See Theoretical Proof). Inference-time scaling-law analyses [45, 31] show that enlarging $n$ drives $R_{\text{ITR}}/n$ towards zero, asymptotically maximising accuracy, and that a perfect verifier ($R_{S}=0$) outperforms self-consistency alone.
Empirically, however, Pass@$k$ diverges from theory if samples are assumed independent (Fig. 5c (left)). Accounting for correlations reveals only $\log n$ independent trajectories; replacing $R_{0}/n$ with $R_{0}/\log n$ restores perfect agreement between prediction and experimental data (Fig. 5c (right)), confirming that the equivalent-circuit framework captures the diminishing returns of modern inference-time scaling-laws.
Figure 6: Optimization Exploration through Magnetic Field Adjustment on ECP.
a, Comparison of accuracy and theoretical power for no‐demo, bottom‐$k$ and top‐$k$ settings, top‐$k$ retrieves for the largest extra voltage. b, Retrieval strategy comparison for random, Diversity-oriented static ICL (DS-ICL) and Similarity-oriented dynamic ICL (SD-ICL) illustrate that SD-ICL achieves the greatest accuracy and power; line plots show both metrics decline as $k$ increases based on Diversity-oriented dynamic ICL (DD-ICL). c, Compared to manual ICL, optimized demonstrations yield significantly higher accuracy and power, with consistent improvement across optimization steps.
3.3 Exploring Novel Optimization Strategies through Magnetic Field Adjustment
When Few-shot ITR Performs Worse than Zero-shot ITR?
LLMs exploit ITL by retrieving task demonstrations, yet the conditions under which demonstrations help remain unclear. We postulated that selecting the top-$k$ most semantically aligned examples would maximise model capacity. To test this idea we compared four retrieval policies: no demonstrations (Demo.), semantically irrelevant (no-meaning) demonstrations, bottom-$k$ and top-$k$ examples. Top-$k$ retrieval improved accuracy across tasks (Fig. 6a). By contrast, no-meaning demonstrations, which exert a negative “semantic field”, reduced effective context utilisation and yielded performance below the zero-shot baseline. These observations show that the magnetic polarity of semantic alignment determines whether demonstrations aid ICL.
The influence of diversity in ITL retrieval: a static vs. dynamic perspective.
Two retrieval paradigms are widely employed: static retrieval, in which a fixed demonstration set is used for all test instances, and dynamic retrieval, in which demonstrations are chosen per instance. Semantic diversity enhances static retrieval by providing complementary contextual signals [23], whereas it can hinder dynamic retrieval by diluting alignment with the query [33]. We quantified these effects with random retrieval, diversity-oriented static retrieval and similarity-oriented dynamic retrieval. Similarity-oriented dynamic retrieval achieved the highest accuracy (57.0%) and power-strength (>70) (Fig. 6b). In static retrieval the same diverse set yielded higher accuracy than random retrieval, but in dynamic retrieval diversity conferred no benefit. Thus, diversity is advantageous only when demonstrations are reused across queries; when selection is instance-specific, alignment dominates.
Demo-Optimization: a novel approach to enhance demonstration efficacy in ITL.
Recently, gradient-based prompt optimisation can refine instructions for large language models [46]. Building on this principle we introduce “demo-optimisation”, an algorithm that jointly adjusts the composition and ordering of demonstrations to balance their relative influence with the user prompt (See Methods in Appendix for more implementation details). As shown in Fig. 6c, demo-optimisation increased accuracy by 20 percentage points relative to instruction-only optimisation and alleviated under-fitting in earlier approaches. The method therefore offers a scalable route to improve demonstration-driven reasoning.
Figure 7: Optimization Exploration through Resistor Reorganization(a, b) and Unified Optimization for Applications (c, d) .